MAGNETIC g_{e}-FACTORS AND ELECTRIC DIPOLE MOMENTS OF LANTHANIDE MONOXIDES: PrO

\underline{HAILING WANG},TIMOTHY C. STEIMLE, Department of Chemistry and Biochemistry, Arizona State University, Tempe,AZ 85287; COLAN LINTON, Physics Department, University of New Brunswick, Fredericton, NB Canada E3B 5A3.

The very complex optical spectra of the lanthanide monoxides are caused by the insensitivity of the electronic energies to the numerous possible arrangements of the Ln^{2+} electrons in the 4f and 6s orbitals. Disentangling the complex optical spectra may be aided by using simple Ligand Field Theory(LFT) to establish the global electronic structure for the lowlying electronic states. A comparison of experimentally determined permanent electric dipole moments, μ_{el}, and magnetic dipole moments, μ_{m}, is an effective means of sorting this myriad of states and assessing the quality of LFT and other electronic structure methodologies. Here we report on the determination of the permanent electric dipole moments, μ_{el}, and magnetic g_{e}-factors for the X_{2}(Ω = 4.5) and [18.1] (Ω = 5.5) states of PrO from the analysis of the optical Stark and Zeeman spectra. The g_{e}-factors are compared with those computed using wavefunctions predicted from ligand field theory^{a}. The μ_{el} value for the X_{2}(Ω = 4.5) state is compared to ab initio^{b}, and density functional^{c} predictions and with the experimental values of other lanthanide monoxides. A phenomenological fit of μ_{el} for the entire series of LnO is used to predict μ_{el} for the isovalent actinide monoxide series.

aCarette, P.,; Hocquet,A.J. Mol. Spectrosc.\underline{131} 301, 1988.

bDolg, M.; Stoll, H. Theor. Chim. Acta. \underline{75},369, 1989.

cWu, Z.; Guan, W. Meng, J. Su, Z. J. Cluster Science \underline{18} 444, 2007.